A firm prints advertising leaflets. They have a set-up fee of £100 and charge 5p per leaflet. Explain in words how to find the total cost of printing N leaflets. Hence explain why the total cost £(C) is given by the formula
C = 100 + 0.05N.
I am sorry to ask for help but I have never been very good with algebra and I am still awaiting a battery for my TI84.

Answer by jim_thompson5910(35256)

No matter how many leaflets you buy, you will pay at least £100. So that's why the constant cost of 100 is in there. Any leaflets you buy, you'll be paying 0.05 (or 5%) of a pound per leaflet. So this is why 0.05N is added.


Here are some examples:

If you buy 100 leaflets, then N=100 and...

C = 100 + 0.05(100) = 100 + 5 = 105

So the total cost is £105



Or if you buy 300 leaflets, then N=300 and...

C = 100 + 0.05(300) = 100 + 15 = 115

So the total cost is £115